The Marginal Revenue function for a company's product is: [05] MR = 220000 - 18x where x equals the number of units sold. If total revenue equals 0 when 0 units are sold, determine the total revenue. 


1
Expert's answer
2021-05-30T23:15:51-0400

Given:

To explain the given statement as follows,

Here, the Marginal Revenue = 220000 - 18x.

To find the total revenue of the company.

Solution:

The Marginal Revenue "\frac{dR}{dx}=220000-18x"

Now integrating with respect to x on both sides, we get

"R(x) = 220000x − 18\frac{x^2}{2}+c"

"R(x) = 220000x\u22129x^2 +c"

"c=0"


Therefore, the Total Revenue as a function of x is 

"R(x) = 220000x\u22129x^2 ."